OpenStudy (anonymous):

write an equation of the line perpendicular to the line given by 2x+3y=7 and having the same y-intercept as the line given by 5x+2y=10.

OpenStudy (anonymous):

First, get these into slope intercept form (y=mx+b) Do you understand how to do that?

OpenStudy (anonymous):

yeah I know how to do that part.

OpenStudy (ghazi):

slope of the required line will be negative inverse of given line that is 3/2 and to find y intercept of the line 5x+2y=10 put x=0 you will get y intercept as 5

OpenStudy (anonymous):

Okay so we get y=-2/3x+7/3 , and y=-5/2x+5

OpenStudy (anonymous):

Here is what I would do: Equation 1: 2x+3y=7 -----> Find the slope, you can do this is if you convert to slope intercept form. Step 2. Use Equation 2. 5x+2y=10.--------> Find the y intercept. If you need more help, just say so.

OpenStudy (anonymous):

Alright, So the equation y=mx+b. Where the "represent the slope, and the "b" represents the y-intecept.

OpenStudy (anonymous):

To be perpendicular, it has to be the inverse slope. So the inverse slope of the line y=-2/3x+7/3 is positive 3/2x . Take the y intercept from the second equation, as required, and plug it into y=mx+b form with our new slope. y=3/2x+5 I believe this is correct.

OpenStudy (anonymous):

The question is asking for the line that is perpendicular to 2x+3y=7, You solved this for the y-interect form, and you got:y=-2/3x+7/3 which is correct. The slope of this line is (-2/3). You want the perpendicular slope of this which is (3/2). So you got that correct.

OpenStudy (anonymous):

thank you all for the help I'm looking over it all now .

OpenStudy (anonymous):

@tumblrfanatic got the correct answer.